OpenStudy (anonymous):

How many bit strings of length 9 contain exactly 5 ones

OpenStudy (dmezzullo):

@Callisto @Preetha

OpenStudy (anonymous):

The answer is \[\left(\begin{matrix}9 \\ 5\end{matrix}\right)\] or\[9C5\]which means the number of ways of selecting 5 things out of 9, called combinations. You can work this out with a calculator or use the formula:\[\left(\begin{matrix}n \\ k\end{matrix}\right) = \frac{ n! }{ (n-k)!(k)! }\]